the length of a rectangle is 3x+2 units and the width is x-6 units, what is the perimeter and area of the rectangle. please help, thx
Answer by stanbon(75887)

the length of a rectangle is 3x+2 units and the width is x-6 units, what is the perimeter and area of the rectangle.
---------------
Perimeter:
P = 2(L+W)
P = 2(3x+2+x-6)
P = 2(4x-4)
P = 8x-8
===================
Area:
A = LW
A = (3x+2)(x-6)
A = 3x^2 -16x -12
====================
Cheers,
Stan H.
=====================